# Scrambled Halton Subsequences and Inverse Star-Discrepancy

### v1

@@ 2 Some Practical Applications
Theorem 2 yields a class of low-discrepancy sequences with three different parameters which can be chosen to minimize the discrepancy. These are the primes, the permutations and the shifts of the scrambled Halton subsequences. They allow to empirically minimize the star-discrepancy for a given $N$ or equivalently minimize the inverse star-discrepancy. We will see that much smaller values than those guaranteed by the bound from Theorem 2 can be realized this way.
@@
To prove the use of our method, we now compare numerical results to related work. To the best of the authors’ knowledge the best empiric results on the inverse star-discrepancy were very recently achieved in [CDP23] by the so-called subset selection algorithm. It provides an heuristic algorithm to find the $m$-element subset $B$ of a bigger set $A\subset[0,1]^{d}$ with $|A|=n$ such that the star-discrepancy is minimal. The algorithm is swap-based and attempts to replace a point of the current subset by one currently not chosen, using the box with the worst local discrepancy.
@@
We compare our results to the values in [CDP23], Table 1, first. We look at dimensions $4$ and $5$ and search for scrambled Halton subsequences which minimize the inverse star-discrepancy. We see in Table 2, that our algorithm yields smaller values for the inverse star-discrepancy than in [CDP23]. To be fair, we stress however that the results from the subset selection algorithm were not explicitly optimized to get particularly small values for the inverse star-discrepancy for all small $N\in\mathbb{N}$ but only some. So there might be some further scope for improvement using this method.
@@
Table 2: Number of points necessary to reach the target discrepancies for Sobol’, subset selection and Scrambled Halton subsequences in dimensions 4 and 5
@@
Since there are by far too many possible combinations of parameters in our method which cannot all be checked, we use the following algorithm to find $d$-dimensional sets with a particularly small star-discrepancy. Motivated by the bounds in Theorem 2, we take the $d$ smallest prime numbers. Then we define a numbers $n_{i}$ of shifts and $m_{i}$ of (randomly chosen) permutations to check for each dimension $i=1,\ldots,d$. Since the number of possible permutations grows exponentially with the dimension, it is reasonable to increase the numbers $m_{i}$ with the $p_{i}$. Due to run time restrictions, $n_{i}$ is therefore defined decreasingly. To further decrease the number of possible permutations, we apply the heuristic from [BW79] and fix $\pi_{i}(0)=0$ for all $i$.
@@
Our algorithm starts with dimension one, where $m_{1}=1$, because there is only one permutation for $p_{1}=2$, and search for the shift factor $a_{1}$ which minimizes the star-discrepancy for the given $N\in\mathbb{N}$. Then $a_{1}$ is fixed and we search for the optimal shift $a_{2}$ (and permutation) such that the two-dimensional star-discrepancy is minimized. As $p_{2}=3$, there are only $2$ permutations with $\pi_{2}(0)=0$ which both can be checked. Next $p_{3}=5$ and the combination of shift $a_{3}$ and permutation $\pi_{3}$ which minimizes the three-dimensional discrepancy (given $a_{1},a_{2},\pi_{1},\pi_{2}$) is searched for. Then we fix $a_{3}$ and $\pi_{3}$, proceed with dimension $4$ and so on.
@@
Table 2 compares the minimal $N$ required to find a set having star-discrepancy below a given threshold when using our algorithm (scrambled Halton), Sobol numbers and the subset selection algorithm. The numbers for the latter two methods are taken from Table 1 in [CDP23]. In dimension $d=4$, the scrambled Halton subsequences require a significantly lower number of points to achieve the listed target discrepancies than the other two methods. In dimension $d=5$, the differences are much smaller but the scrambled Halton subsequences are still superior. Note that even small changes can be relevant in practice, where one simulation (based on the assumptions defined by the point in the set) can take several hours and thus produce high computational and economic costs.
Moreover, our algorithm does not only yield smaller values of $N$ but has two other benefits in comparison to the subset selection algorithm: another main advantage of our algorithm is that we can explicitly write down the parameters for the best found scrambled Halton subsequences, so that the reader can independently reproduce our results. We report them for the results in Table 2 in Table 4 and Table 5 of the Appendix 4. Also the running time of our algorithm is rather low and not very demanding in terms of programming skills given we have at hand a fast algorithm to calculate the star-discrepancy.
@@
Another approach to find sets with a particularly small discrepancy for a given $N$ was introduced in [DR13]. The main components of the algorithm therein are based on evolutionary principles. The algorithm is called optimized Halton and was the first algorithm which could be adapted easily to optimize the inverse star-discrepancy. Therefore, it can serve as another good point of comparison. For completeness, we also add to it the results from [DGW10], which served as a benchmark in [DR13].
@@
In Table 3, we see that the scrambled Halton subsequences approach yields promissing values for the best found star-discrepancy given the dimension $d$ and the set size $N$. While the discrepancy of the scrambled Halton subsequences is the smallest for dimensions $d=5$ and $d=7$, it is slightly worse in the case $d=9$. Since the results are not deterministic but depend on the random choices of the permutations, the two approaches here and in [DR13] should in our opinion therefore be regarded as equally competitive. The results from [DGW10] on the other hand yield much larger values for the discrepancy of the best found set.
@@
Table 3: Exact discrepancy results for the sequences presented in [DR13] and [DGW10].
@@ 3 Theoretical background
In this section we present the proofs of our theoretical results. We start by showing Theorems 1 and 2. For this purpose we introduce concepts from p-adic discrepancy theory first, see e.g. [Mei68, Wei24] for details.
@@
Recall that a polynomial $f:\mathbb{Z}\to\mathbb{Z}$ is called a permutation polynomial $\textrm{mod}\ n$ for $n\in\mathbb{N}$ if it is bijection on $\mathbb{Z}/n\mathbb{Z}$ and thus a permutation of the elements in $\mathbb{Z}/n\mathbb{Z}$. It turns out to be useful to regard $f(n)$ as a sequence in the $p$-adic integers $\mathbb{Z}_{p}=\{x\in\mathbb{Q}_{p}\,:\,|x|_{p}\leq 1\}$, where $\left|\cdot\right|_{p}$ denotes the $p$-adic absolute value and $\mathbb{Q}_{p}$ is the completion of $\mathbb{Q}$ with respect to it. For a sequence $(x_{n})\subset\mathbb{Z}_{p}$ and $N\in\mathbb{N}$, the $p$-adic discrepancy is defined as
$\delta_{N}^{p}(x_{n}):=\sup_{z\in\mathbb{Z}_{p},k\in\mathbb{N}}\left|\frac{\#\left(\textrm{Disc}_{p}(z,k)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\frac{1}{p^{k}}\right|$
where the p-adic disc at center $z\in\mathbb{Z}_{p}$ of radius $p^{-k}$ is given by
@@
By the same argument as in the real setting, it is straightforward to see that $\tfrac{1}{N}\leq D_{N}(x_{n})\leq 1$ holds for all sequences $(x_{n})\subset\mathbb{Z}_{p}$ and all $N\in\mathbb{N}$. The proof of Theorem 1 now essentially follows the arguments in [Wei24].
Proof of Theorem 1.
If $f:\mathbb{N}\to\mathbb{N}$ is a permutation polynomial $\bmod\ p^{2}$, it is a permutation polynomial $\bmod\ p$ as well. By Hensel’s Lemma it is also a permutation polynomial $\bmod\ p^{k}$ for all $k\in\mathbb{N}$ with $k\geq 3$.
As $f$ is a permutational polynomial mod $p^{k}$ for all $k$, every disk $\textrm{Disc}_{p}(z,k)$ contains $\lfloor Np^{-k}\rfloor$ or $\lfloor Np^{-k}\rfloor+1$ elements. This property remains true after applying the permutation $\pi$ which only permutes the residue classes $\bmod\ p^{k}$. Hence, the $p$-adic discrepancy satisfies $\delta_{N}^{p}(f(n))\leq\frac{1}{N}$ for all $N\in\mathbb{N}$. The following theorem by Meijer in [Mei68] thus implies the claim after realizing that $\varphi_{b}(f(n))$ therein equals the definition of the scrambled van der Corput sequene in (1).
Theorem 5.
Let $(x_{n})\subset\mathbb{Z}_{p}$ be an arbitrary sequence and denote its p-adic discrepancy for $N\in\mathbb{N}$ by $\delta_{N}^{p}$. Let $d_{N}$ be the discrepancy of the corresponding (real) sequence $(\varphi_{p}(x_{n}))\subset[0,1)$, where
@@
$\delta_{N}^{p}<d_{N}<\delta_{N}^{p}\left(2+\frac{2(p-1)}{\log p}\log\left(\delta_{N}^{-1}\right)\right).$
@@
$\mathbb{Z}_{P}:=\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{1}}\times\ldots\times\mathbb{Z}_{p_{d}}.$
@@
in $\mathbb{Z}_{P}$. The normalized Haar measure on $\mathbb{Z}_{P}$ is denoted by $\mu$ such that we obtain $\mu(\textrm{Disc}_{P}(Z,K))=\prod_{i=1}^{d}p_{i}^{-k_{i}}$. For a sequence $(x_{n})$ in $\mathbb{Z}_{P}$ its discrepancy is then defined by
$\delta_{N}^{P}(x_{n}):=\sup_{Z\in\mathbb{Z}_{P},k\in\mathbb{N}_{0}^{d}}\left|\frac{\#\left(\textrm{Disc}_{P}(Z,K)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\mu\left(\textrm{Disc}_{P}(Z,K)\right)\right|.$
@@
Theorem 6.
Let $P=(p_{1},\ldots,p_{d})$ be a vector of distinct primes and let $(x_{n})$ be a sequence in $\mathbb{Z}_{P}$. If $\delta_{N}^{P}$ denotes the $P$-adic discrepancy of $(x_{n})$ in $\mathbb{Z}_{P}$ and $d_{N}$ the discrepancy of $\varphi_{P}(x_{n})$ in $\mathbb{R}^{d}$, then it holds that
$2^{-d}\delta_{N}^{P}\leq d_{N}\leq\delta_{N}^{P}\left(2\sum_{k=1}^{d}g_{d}+\log\left(\left(\delta_{N}^{p}\right)^{-1}\right)^{d}\prod_{k=1}^{d}\frac{2(p_{k}-1)}{\log(p_{k})}\right).$
In order to show Theorem 2 it thus suffices to prove that $\delta_{N}^{P}(x_{n})=\frac{1}{N}$ for scrambled Halton subsequences.
@@
Since $f_{i}(n)$ is a permutation polynomial $\bmod\ p_{i}^{k_{i}}$ for all $i=1,\ldots,d$ and all $k_{i}\in\mathbb{N}_{0}$ and $\textrm{Disc}_{P}(Z,K)$ was arbitrary, this property regarding the number of elements remains true for the sequence $f(n)=(f_{1}(n),\ldots,f_{d}(n))$. Finally, also the $\pi_{i}$ only permute the residue classes $\bmod\ p_{i}^{k}$ of the sequence elements which does neither affect the fact that the number of elements in $\textrm{Disc}_{P}(Z,K)$ is $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor$ or $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor+1$. Thus the discrepancy is $\delta_{N}^{P}\leq\frac{1}{N}$.
To see that the discrepancy is not equal to 00 choose an arbitrary point $Z=(z_{1},\ldots,z_{d})\in\mathbb{Z}_{P}$ which is an element of the sequence. Then take $k_{1}$ large enough such that $\textrm{Disc}_{p_{1}}(z_{1},k_{1})$ does not contain any other element of the sequence. Thus $\delta_{N}^{P}\geq\frac{1}{N}$ and the claim follows. ∎
We will not now anymore apply p-adic discrepancy theory but proceed with the proof of Theorem 4. For small dimensions we will use a standard construction based on Halton sequences to ensure that the bound $D_{N}^{*}(x_{n})\leq 2.4631387\sqrt{\frac{d}{N}}$ is satisfied. These are the so-called Hammersley point sets which lift Halton sequences in dimension $d$ to sets in dimension $d+1$. Furthermore, our proof relies on the concept of bracketing numbers which we introduce afterwards as well.

### v2

@@ 2 Some Practical Applications
Theorem 2 yields a class of low-discrepancy sequences with three different parameters which can be chosen to minimize the star-discrepancy. These are the primes, the permutations and the shifts of the scrambled Halton subsequences. They allow to empirically minimize the star-discrepancy for a given $N$ or equivalently minimize the inverse star-discrepancy. We will see that even much smaller values than $\frac{\log(N)^{d}}{N}\prod\frac{p_{i}-1}{2\log(p_{i})}$, which would be the best we can hope for according to Theorem 2 when ignoring the $O(\tfrac{\log(N)^{d-1}}{N})$ term, can be realized this way.
One-dimensional sequences.
At first, we again discuss the one-dimensional situation. For a sound comparison of our approach with the original and the scrambled van der Corput sequence it should be born in mind that the star-discrepancy is by our method minimized for a given $N$ which might mean that it is particularly small for the chosen $N$ but not for many other $n<N$. In Figure 1, we therefore plot the star-discrepancy values of the usual van der Corput sequence $\varphi_{3}$, the scrambled van der Corput sequence $\varphi_{3,\pi}$ according to [BW79] and the best scrambled van der Corput subsequence $\varphi_{3,\pi,a}$ for $N=1,000$ according to our approach for $0001\leq n\leq 1,$ 000. The values for the star-discrepancy are rescaled by $n/\log(n)$ so that the plots remain readable throughout and the asymptotic behavior is better visible. In Figure 2, the same is done for $p=5$ instead. The scrambled sequence as well as the scrambled subsequence almost systematically outperform the original van der Corput sequence, while the former two methods have the lowest star-discrepancy for approximately the same number of $n\leq 1,000$.
An additional interesting question concerns how well the sequences $\varphi_{p,\pi,a}$ perform beyond the anchoring value $N$. In Figure 4 of the appendix, we expanded Figure 2 up to $N=5,000$ and the empirical observations remain the same. The sequences $\varphi_{\pi,5}$ and $\varphi_{\pi,5,a}$ have the lowest star-discrepancy for about the same number of $n$ and both clearly outperform the original van der Corput sequence. Similar observations can be made for other values of $p$.
In a more comprehensive setting, the star-discrepancy of scrambled van der Corput sequences may also be compared to other low-discrepancy sequences. We would like to address here two specific examples which both are known to have a particularly good asymptotic behavior, see e.g. [Clé24]. The first one is the golden ratio Kronecker sequence defined by $(x_{n})=\{n\phi\}$, where $\phi=\frac{1+\sqrt{5}}{2}$ is the golden ratio. The other, more recent one, is the Kritzinger-sequence $Kri_{n}$ introduced in [Kri21]: set $Kri_{1}=\frac{1}{2}$ and define $Kri_{n+1}$ for $n\geq 1$ by
$Kri_{n+1}=\arg\min_{x\in[0,1)]}-2\sum_{k=1}^{n}\max(Kri_{k},x)+(x+1)x^{2}-x.$
If there are several solutions, then always the smallest one is chosen as $Kri_{n+1}$. It can then be shown that all elements are of the form $Kri_{n}=\frac{2k-1}{2n}$ for some $1\leq k\leq n$.
In Figure 3, we compare the rescaled star-discrepancy of these two sequences with $\varphi_{5,\pi,a}(1000)$ and $\varphi_{5,\pi,a}(2000)$ up to $n=2,000$, where the number in the bracket indicates the anchoring value $N$ of the scrambled van der Corput subsequences. It turns out that $\varphi_{5,\pi,a}(1000)$ has the lowest star-discrepancy for $150$ different values of $n$ and $\varphi_{5,\pi,a}(2000)$ for $475$ different values while the best value is achieved for the golden ratio Kronecker sequence in $637$ cases and for the Kritzinger sequence in $737$. In a direct comparison of $\varphi_{5,\pi,a}(2000)$ and the Kritzinger sequence, $Kri_{n}$ wins $1,250$ times. However, it is not surprising that $\varphi_{5,\pi,a}(2000)$ is the winner of $80$ of the last $100$ values of $n$. Similarly, $\varphi_{5,\pi,a}(1000)$ is the winner of all four sequences $41$ times in the range $n=901,\ldots,1.100$. Summing up, scrambled van der Corput sequences seem to be good candidates when searching for sequences with a particularly small local star-discrepancy in a given (small) range but they also yield good star-discrepancy values for other $n$.
Figure 3: Comparison of $\varphi_{5},\varphi_{5,\pi}$ and $\varphi_{5,\pi,1,000}$ up to $N=2,000$.
Multi-dimensional setting.
@@
To further prove the use of our method, we compare more numerical results to related work in the multi-dimensional case. To the best of the authors’ knowledge the best empiric results on the inverse star-discrepancy were very recently achieved in [CDP23] by the so-called subset selection algorithm. It provides an heuristic algorithm to find the $m$-element subset $B$ of a larger set $A\subset[0,1]^{d}$ with $|A|=n$ such that the star-discrepancy is minimal. The algorithm is swap-based and attempts to replace a point of the current subset by one currently not chosen, using the box with the worst local discrepancy.
@@
We compare our results to the values in [CDP23], Table 1, first. We look at dimensions $4$ and $5$ and search for scrambled Halton subsequences which minimize the inverse star-discrepancy. We see in Table 2, that our algorithm yields smaller values for the inverse star-discrepancy than in [CDP23]. To be fair, we stress, however, that the results from the subset selection algorithm were not explicitly optimized to get particularly small values for the inverse star-discrepancy for all small $N\in\mathbb{N}$ but only some specific values $N$. So there might be some further scope for improvement using this method.
@@
Table 2: Number of points necessary to reach the target star-discrepancies for Sobol’, subset selection and Scrambled Halton subsequences in dimensions 4 and 5
@@
Since there are by far too many possible combinations of parameters in our method which cannot all be checked, we use the following algorithm to find $d$-dimensional sets with a particularly small star-discrepancy. Motivated by the bounds in Theorem 2, we take the $d$ smallest prime numbers. Then we define a numbers $n_{i}$ of shifts and $m_{i}$ of (randomly chosen) permutations to check for each dimension $i=1,\ldots,d$. Since the number of possible permutations grows faster than exponentially with the dimension, it is reasonable to increase the numbers $m_{i}$ with the $p_{i}$. Due to run time restrictions, $n_{i}$, i.e. the maximum value for the shift, is therefore defined decreasingly. To further decrease the number of possible permutations, we apply the heuristic from [BW79] and restrict ourselves to permutations $\pi_{i}$ such that $\pi_{i}(0)=0$ for all $i$.
@@
For fixed $N\in\mathbb{N}$, our algorithm starts with dimension one, where $m_{1}=1$, because there is only one permutation for $p_{1}=2$, and search for the shift factor $1\leq a_{1}\leq n_{1}$ which minimizes the one-dimensional star-discrepancy. Then $a_{1}$ is fixed and we search for the optimal shift in the range $1\leq a_{2}\leq n_{2}$ (and permutation) such that the two-dimensional star-discrepancy is minimized. As $p_{2}=3$, there are only $2$ permutations with $\pi_{2}(0)=0$ which both can be checked. Next $p_{3}=5$ and the combination of shift $a_{3}$ and permutation $\pi_{3}$ which minimizes the three-dimensional star-discrepancy (given $a_{1},a_{2},\pi_{1},\pi_{2}$) is searched for. Then we fix $a_{3}$ and $\pi_{3}$, proceed with dimension $4$ and so on.
@@
Table 2 compares the minimal $N$ required to find a set having star-discrepancy below a given threshold when using our algorithm (scrambled Halton), Sobol numbers and the subset selection algorithm. The numbers for the latter two methods are taken from Table 1 in [CDP23]. In dimension $d=4$, the scrambled Halton subsequences require a significantly lower number of points to achieve the listed target star-discrepancies than the other two methods. In dimension $d=5$, the differences are much smaller but the scrambled Halton subsequences are still superior. Note that even small changes can be relevant in practice, where one simulation (based on the assumptions defined by the point in the set) can take several hours and thus produce high computational and economic costs.
Moreover, our algorithm does not only yield smaller values of $N$, but has two other benefits in comparison to the subset selection algorithm: first, we can explicitly write down the parameters for the best found scrambled Halton subsequences in a compact form, so that the reader can independently reproduce our results. We report them for the results in Table 2 in Table 5 and Table 6 of the Appendix 4. Of course, it is also possible to write down the parametrization, i.e. the random seed, of the subset selection algorithm. However, in our eyes, this would be a less self-contained representation of the sets. Second, Halton subsequences are sequences and therefore after optimization for a given $N$, they can be extended to an arbitrary large $N_{1}>0$ by just using the given parametrization. We have seen in Figure 3 for the case of dimension $1$ that these sequences typically perform very well if $N_{1}$ is not too far apart from $N$.
@@
Another approach to find sets with a particularly small star-discrepancy for a given $N$ was introduced in [DR13]. The main components of the algorithm therein are based on evolutionary principles. The algorithm is called optimized Halton and was the first algorithm which could be adapted easily to optimize the inverse star-discrepancy. Therefore, it can serve as another good point of comparison. For completeness, we also add to it the results from [DGW10], which served as a benchmark in [DR13].
@@
In Table 3, we see that the scrambled Halton subsequences approach yields promising values for the best found star-discrepancy given the dimension $d$ and the set size $N$. While the star-discrepancy of the scrambled Halton subsequences is the smallest for dimensions $d=5$ and $d=7$, it is slightly worse in the case $d=9$. Since the results are not deterministic but depend on the random choices of the permutations, the two approaches here and in [DR13] should in our opinion therefore be regarded as equally competitive. The results from [DGW10] on the other hand yield much larger values for the star-discrepancy of the best found set.
@@
Table 3: Exact star-discrepancy results for the sequences from [DR13] and [DGW10] and scrambled Halton sequences.
Two Dimensions.
Two dimensions may be regarded as an intermediate case when using our approach: on the one hand, it can be treated as a two-dimensional minimization task for the star-discrepancy of Halton sequences. On the other hand, it may also be seen as one-dimensional minimization task for the (two-dimensional) scrambled Hammersley point set defined by $y_{n}=(n/N,x_{n})\in[0,1)^{2}$ for $n=1,\ldots,N-1$, where $(x_{n})$ is a one-dimensional scrambled van der Corput subsequence. We will discuss the theoretical properties of Hammersley point sets in more detail in Section 3.
Also the case of two dimensions has lately attracted particular attention: it has been treated with the machine learning approach fro, [RKB+24], which we have already mentioned. In [CDKP24], two-dimensional sets are constructed by considering two lists of coordinates and an assignment matrix, linking the list of coordinates to the points’ position. This way, the optimization of the star-discrepancy can also be formulated as a problem of finding an optimal permutation. The base coordinates considered therein are $y_{n}=(n/N,x_{n})\in[0,1)^{2}$ for $n=1,\ldots,N-1$, where $(x_{n})$ is a low-discrepancy sequence, as well. In particular the Kritzinger sequence and the Kronecker sequence for the golden ratio are taken for $(x_{n})$. In the latter case, the set is also called Fibonacci point set.
In our setting, it turned out that the Hammersley point set yields much smaller discrepancy values than two-dimensional Halton sequences. Therefore, the presentation in this article restricts itself to Hammersley point sets. To be precise, we considered the Hammersley point set for the $9$ smallest primes and the allowed number of shifts were $1,000,500$ and $100$ for the first three prime numbers and $50$ for all others. In order to give every prime the same chance to be the winner the numbers of random permutations are set such that the total number of tries is always $1,000$.
În Table 4, we compare our approach with the best results obtained by the two alternative methods. It turns out that the star-discrepancy of the scrambled Hammersley point set is smaller than for the Fibonacci point set. Nonetheless, it is outperformed by both alternative methods. Still it might be promising to use scrambled Hammersley point sets as base coordinates in the approach of [CDKP24]. Moreover, we see from Table 7 in the appendix, that the configuration $p=2$ and the shift $n=509$ is always the best found scrambled Hammersley point set for $n\geq 260$ and should be a particularly good candidate.
$N$ Fibonacci Scrambled Ham. Result [RKB+24] Best result from [CDKP24]
$20$ $0.105833$ $0.0800$ $0.00666$ $0.06219$
$50$ $0.049165$ $0.0395$ $0.2742$
$100$ $0.0026105$ $0.0200$ $0.0188$ $0.01492$
$180$ $0.015165$ $0.0131$ $0.0115$ $0.00901$
$220$ $0.012407$ $0.0100$ $0.0097$ $0.00737$
$260$ $0.010894$ $0.0098$ $0.0084$ $0.00640$
$350$ $0.008731$ $0.0075$ $0.004872$
$420$ $0.00728$ $0.0065$ $0.0058$ $0.00412$
$500$ $0.00611$ $0.0055$ $0.0052$
Table 4: Comparison of star-discrepancy for scrambled Hammersley point sets and results from [RKB+24] and [CDKP24] as reported in [CDKP24].
More related work.
Finally, we do not want to sweep under the rug that other choices for the permutations $\pi$ than those in [BW79] have been made by several authors, both in the one- and multidimensional. The excellent article [FL09] contains a comprehensive overview of many of these approaches, as well as numerical results that evaluate their performance in certain integration tasks. However, all of these methods are purely based on optimizing the permutation and do not include any shifts. Since all of these methods date prior to 2009, we decided to use more recent results as point of comparison in this article.
@@ 3 Theoretical background
In this section we present the proofs of our theoretical results. We start by showing Theorems 1 and 2. For this purpose we introduce concepts from $p$-adic discrepancy theory first, see e.g. [Mei68, Som22, Wei25] for details.
@@
Recall that the $p$-adic absolute value $|\cdot|_{p}$ is for $a=\frac{b}{c}$ with $b,c\in\mathbb{Z}\setminus\{0\}$ defined as $|a|_{p}:=p^{-m}$, where $m$ is the highest possible power with $a=p^{m}\frac{b^{\prime}}{c^{\prime}}$ and $(b^{\prime}c^{\prime},p)=1$. It turns out to be useful to regard the values $f(n)$ of a permutation polynomial $f:\mathbb{N}\to\mathbb{N}$ as a sequence in the $p$-adic integers $\mathbb{Z}_{p}=\{x\in\mathbb{Q}_{p}\,:\,|x|_{p}\leq 1\}$, where $\mathbb{Q}_{p}$ are the $p$-adic numbers, i.e. the completion of $\mathbb{Q}$ with respect to $|\cdot|_{p}$. For a sequence $(x_{n})\subset\mathbb{Z}_{p}$ and $N\in\mathbb{N}$, the $p$-adic discrepancy is defined as
$\delta_{N}^{(p)}(x_{n}):=\sup_{z\in\mathbb{Z}_{p},k\in\mathbb{N}}\left|\frac{\#\left(\textrm{Disc}_{p}(z,k)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\frac{1}{p^{k}}\right|$
and the $p$-adic disc at center $z\in\mathbb{Z}_{p}$ of radius $p^{-k}$ is given by
@@
The concept of $p$-adic discrepancy allows to quantify the degree of uniformity in $\mathbb{Z}_{p}$ in a similar way as real discrepancy theory does in $[0,1]$. The notion originally stems from [Cug62]. By the same argument as in the real setting, it is straightforward to see that $\tfrac{1}{N}\leq D_{N}(x_{n})\leq 1$ holds for all sequences $(x_{n})\subset\mathbb{Z}_{p}$ and all $N\in\mathbb{N}$. However, the lower bound is in fact also sharp for sequences. Unlike in the real case there is no $\log(N)$ term required in the numerator as was shown in [Bee69].
The proof of Theorem 1 now essentially follows the arguments in [Wei25] and relies on Hensel’s Lemma.
Lemma 5 (Hensel’s Lemma).
Let $f(x)\in\mathbb{Z}[X]$ and let $a\in\mathbb{Z}$ with $f(a)\equiv 0\bmod\ p^{k},$ but $f^{\prime}(a)\not\equiv\bmod\ p^{k}$. Then there exist $b\in\mathbb{Z}$ with $f(b)\equiv 0\bmod\ p^{k+1}$ and $a\equiv b\bmod\ p^{k}$.
Another key ingredient is the following theorem of Meijer.
Theorem 6 (Meijer, [Mei68]).
Let $(x_{n})\subset\mathbb{Z}_{p}$ be an arbitrary sequence and denote its $p$-adic discrepancy for $N\in\mathbb{N}$ by $\delta_{N}^{(p)}$. Let $D_{N}$ be the extreme discrepancy of the corresponding (real) sequence $(\varphi_{p}(x_{n}))\subset[0,1)$, where
@@
$\delta_{N}^{(p)}<D_{N}<\delta_{N}^{(p)}\left(2+\frac{2(p-1)}{\log p}\log\left((\delta_{N}^{(p)})^{-1}\right)\right).$
Proof of Theorem 1.
If $f:\mathbb{N}\to\mathbb{N}$ is a permutation polynomial $\bmod\ p^{2}$, it is a permutation polynomial $\bmod\ p$ as well and there are no solutions to $f^{\prime}(x)\equiv 0\mod p$. Now let $y\in\mathbb{Z}/p^{k}\mathbb{Z}$ and denote by $\overline{f}$ the reduction of $f\mod p$ and by $\overline{y}$ the reduction of $y\mod p$. By assumption there is a simple root for the polynomial $\overline{f}(x)-\overline{y}\equiv 0\mod p$, so Hensel’s lemma inductively implies that there is a unique solution to $f(x)\equiv y\mod p^{k}$. In other words a solution of the equation $\overline{f}(x)-\overline{y}\equiv 0\mod p$ can be lifted to a solution $\bmod\ p^{k}$ for arbitrary $k\in\mathbb{N}$ by Hensel’s lemma.
As $f$ is a permutational polynomial mod $p^{k}$ for all $k$, every disk $\textrm{Disc}_{p}(z,k)$ contains $\lfloor Np^{-k}\rfloor$ or $\lfloor Np^{-k}\rfloor+1$ elements of the sequence. This property remains true after applying the permutation $\pi$ which only permutes the residue classes $\bmod\ p^{k}$. Hence, the $p$-adic discrepancy satisfies $\delta_{N}^{(p)}(f(n))\leq\frac{1}{N}$ for all $N\in\mathbb{N}$. Theorem 6 thus implies the claim after realizing that $\varphi_{b}(f(n))$ therein equals the definition of the scrambled van der Corput sequence in (1). ∎
@@
$\mathbb{Z}_{P}:=\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\ldots\times\mathbb{Z}_{p_{d}}.$
@@
in $\mathbb{Z}_{P}$. The normalized Haar measure on $\mathbb{Z}_{P}$ is denoted by $\mu$ such that we obtain $\mu(\textrm{Disc}_{P}(Z,K))=\prod_{i=1}^{d}p_{i}^{-k_{i}}$. For a sequence $(x_{n})$ in $\mathbb{Z}_{P}$ its $p$-adic discrepancy is then defined by
$\delta_{N}^{(P)}(x_{n}):=\sup_{Z\in\mathbb{Z}_{P},k\in\mathbb{N}_{0}^{d}}\left|\frac{\#\left(\textrm{Disc}_{P}(Z,K)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\mu\left(\textrm{Disc}_{P}(Z,K)\right)\right|.$
@@
Theorem 7 (Meijer, [Mei68]).
Let $P=(p_{1},\ldots,p_{d})$ be a vector of distinct primes and let $(x_{n})$ be a sequence in $\mathbb{Z}_{P}$. If $\delta_{N}^{P}$ denotes the $P$-adic discrepancy of $(x_{n})$ in $\mathbb{Z}_{P}$ and $D_{N}$ the extreme discrepancy of $\varphi_{P}(x_{n})$ in $\mathbb{R}^{d}$, then it holds that
$2^{-d}\delta_{N}^{(P)}\leq D_{N}\leq\delta_{N}^{(P)}\left(2\sum_{k=1}^{d}g_{d}+\log\left(\left(\delta_{N}^{(P)}\right)^{-1}\right)^{d}\prod_{k=1}^{d}\frac{2(p_{k}-1)}{\log(p_{k})}\right).$
In order to show Theorem 2 it thus suffices to prove that $\delta_{N}^{(P)}(x_{n})=\frac{1}{N}$ for scrambled Halton subsequences.
@@
Since $f_{i}(n)$ is a permutation polynomial $\bmod\ p_{i}^{k_{i}}$ for all $i=1,\ldots,d$ and all $k_{i}\in\mathbb{N}_{0}$ and $\textrm{Disc}_{P}(Z,K)$ was arbitrary, this property regarding the number of elements remains true for the sequence $f(n)=(f_{1}(n),\ldots,f_{d}(n))$. Finally, also the $\pi_{i}$ only permute the residue classes $\bmod\ p_{i}^{k}$ of the sequence elements which does neither affect the fact that the number of elements in $\textrm{Disc}_{P}(Z,K)$ is $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor$ or $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor+1$. Thus, the $p$-adic discrepancy is $\delta_{N}^{(P)}\leq\frac{1}{N}$.
To see that the $p$-adic discrepancy is not equal to 00 choose an arbitrary point $Z=(z_{1},\ldots,z_{d})\in\mathbb{Z}_{P}$ which is an element of the sequence. Then take $k_{1}$ large enough such that $\textrm{Disc}_{p_{1}}(z_{1},k_{1})$ does not contain any other element of the sequence. Thus $\delta_{N}^{(P)}\geq\frac{1}{N}$ and the claim follows. ∎
Now, we will no longer apply the $p$adic discrepancy theory, but proceed with the proof of Theorem 4. For small dimensions we will use a standard construction based on Halton sequences to ensure that the bound $D_{N}^{*}(x_{n})\leq 2.4631832\sqrt{\frac{d}{N}}$ is satisfied. These are the so-called Hammersley point sets which lift Halton sequences in dimension $d$ to sets in dimension $d+1$. Furthermore, our proof relies on the concept of bracketing numbers which we introduce afterwards as well.